The pressure of an ideal gas is reduced by 50%, resulting in a decrease in temperature to 75% of the initial value. Calculate the ratio of the final to initial volumes of the gas. 


1
Expert's answer
2021-06-21T11:39:08-0400

Initial pressure: "P_0"

Final pressure: "P_1=0.5P_0"

Initial temperature: "T_0"

Final temperature: "T_1=1.75T_0"

According to the combined gas law:


"\frac{P_0V_0}{T_0}=\frac{P_1V_1}{T_1},\\\space\\
\frac{V_1}{V_0}=\frac{T_1P_0}{T_0P_1}=\frac{1.75·1}{1·0.5}=3.5."